OpenStudy (anonymous):

find all roots 5x^2+5x+60=0

OpenStudy (luigi0210):

Hello there, welcome to Openstudy :) First start off by factoring out a 5

OpenStudy (wolf1728):

Basically, divide the equation by 5 x² +x +12 =0

OpenStudy (anonymous):

what happens after x^2+x+12=0

OpenStudy (wolf1728):

Then you factor the equation. You know how to do that?

OpenStudy (anonymous):

quadratically? i dnt know i get stuck around here

OpenStudy (wolf1728):

Hmm I don't think that factors easily. Want to try the quadratic formula?

OpenStudy (anonymous):

ya but is there another formula?

OpenStudy (wolf1728):

no that's the only one

OpenStudy (anonymous):

ok

OpenStudy (wolf1728):

we can solve it before it is divided by 5 or after

OpenStudy (anonymous):

can you show me how to solve it before its divided by five

OpenStudy (wolf1728):

Sure. The equation is 5x²+5x+60=0 a = 5 b = 5 c = 60

OpenStudy (wolf1728):

quadratic formula is x = [-b +-square root (b² - 4*a*c)] / 2a

OpenStudy (wolf1728):

Can you fill in the formula with the a b and c values?

OpenStudy (anonymous):

ya im doin it now

OpenStudy (wolf1728):

Okay, then I'll wait for you to finish.

OpenStudy (anonymous):

would it be -5+-square root(-1150)

OpenStudy (anonymous):

i think its wrong

OpenStudy (wolf1728):

Well I'll fill in the numbers hee: a = 5 b=5 c = 60 x = [-5 +-square root(25 -4*5*60)] / 10 x= [-5 +-square root(25 -1200)] / 10 x= [-5 +-square root(-1175)] / 10

OpenStudy (wolf1728):

Seems as if you didn't have the divided by 2a in there.

OpenStudy (anonymous):

ya i forgot to write it and can it be simplified?

OpenStudy (wolf1728):

Sure but as you can see, if you remembered the 2a then it would have been filled in correctly.

OpenStudy (wolf1728):

No, it should be square root of 1175 and not 1150

OpenStudy (anonymous):

ya i messed up on that

OpenStudy (wolf1728):

Well we can see about simplifying this. x= [-5 +-square root(-1175)] / 10 x = -5/10 +-square root(-47*25) / 10 x = -1/2 (+- 5* square root(-47) / 10) x = -1/2 (+- 1/2* square root(-47))

OpenStudy (anonymous):

it cnt be simplified anymore?

OpenStudy (anonymous):

what would be the roots of the equation?

OpenStudy (wolf1728):

You could neaten it up by posting actual values. x = -1/2 (+- 1/2* square root(-47)) square root of -47 is 6.8556546004 * i so x = -1/2 + 3.4278273002 * i and x = -1/2 - 3.4278273002 * i

OpenStudy (wolf1728):

yes - it's just that easy LOL

OpenStudy (anonymous):

lmao thanks so the roots would be -1/2-3.4278273002

OpenStudy (anonymous):

?

OpenStudy (wolf1728):

You forgot to add the 'i' at the end. ('i' is the square root of minus 1) -1/2 + 3.4278273002 * i and -1/2 - 3.4278273002 * i

OpenStudy (anonymous):

i didnt know that why did u add the *

OpenStudy (wolf1728):

just to make the 'i' stand out a little more because it would be easy to mistake the i for a 1

OpenStudy (anonymous):

o ok thanks a lot

OpenStudy (wolf1728):

well that's it - I trust you can use the formula a little better now.

OpenStudy (anonymous):

def.

OpenStudy (wolf1728):

okay - see you later

OpenStudy (anonymous):

cya